Please fill in all the blanks to the below table and give reasoning behind the answers.

Address Class

Range

Address Bytes

Number of

Networks

Host Bytes

Number of

Hosts

Class A

1 -126

126

16,777,214

Class B

2

2

Class C

192-223
Answer #1

Address Class

Range

Address Bytes

Number of

Networks

Host Bytes

Number of

Hosts

Class A

1 -126

1

128-2=126

3

16,777,214

Class B

127-191

2

16384-2=16380

2

65,536

Class C

192-223

3 2097152-2=2097150 1 256